OpenStudy (anonymous):

544x=425eight find the base x

OpenStudy (anonymous):

Thanks Ranga for helping me out again

OpenStudy (ranga):

First convert 425 base 8 to decimal Remember the way we converted various base numbers to the decimal system before. Do the same here. To convert a three digit number say ABC base 8 to decimal, we do the following: C + B(8^1) + A(8^2) Can you convert 425 base 8 to decimal and tell me what you get?

OpenStudy (anonymous):

ok give me a sec here

OpenStudy (anonymous):

651?

OpenStudy (anonymous):

I think my brain is fried been doing these problems for 3 days

OpenStudy (ranga):

Try again. Let me show it clearly by using the Equation button below the reply box. Give me a second.

OpenStudy (anonymous):

ok

OpenStudy (ranga):

\[425_{8} = 5 + (2)(8^{1}) + (4)8^{2} = ?\]

OpenStudy (anonymous):

now do i times the 8 x2 or add them?

OpenStudy (anonymous):

83?

OpenStudy (ranga):

you do 2 times 8 then you do 4 times 8 times 8 then you add all three numbers

OpenStudy (anonymous):

ok just sec

OpenStudy (anonymous):

267

OpenStudy (ranga):

close but not quite

OpenStudy (anonymous):

did I just do my math wrong? did it 3 times now?

OpenStudy (ranga):

2 x 8 = ?

OpenStudy (anonymous):

16

OpenStudy (ranga):

ok. 4 x 8 x 8 = ?

OpenStudy (anonymous):

256 which I got those 2

OpenStudy (ranga):

5 + 16 + 256 = ?

OpenStudy (anonymous):

OH I see what I did duh....277

OpenStudy (ranga):

OK. Now on the left you have 544 base x. Don't let the x throw you off. Follow the same rule as before to convert it to decimal.

OpenStudy (ranga):

\[544_{x} = 4 + (4)(x ^{1}) + (5)(x ^{2}) = 5x ^{2} + 4x + 4\]

OpenStudy (anonymous):

ok just sec

OpenStudy (ranga):

I have given you the answer above about how to convert base x to decimal. Go through it and see if it is clear or ask me if you have doubts.

OpenStudy (anonymous):

yes I do have doubts about adding the x's together I know dont let it throw me off but it does

OpenStudy (anonymous):

5x(little2) =4x=4 does throw me off

OpenStudy (anonymous):

I meant + where there is =

OpenStudy (ranga):

544 base x The last term 4 stays as 4. The next term to the left is also a 4 but it has to be multiplied by x (which is the base) raised to 1. So it will be 4x^1 = 4x The next term to the left is 5 and that has to be multiplied by x^2 so that will be 5x^2 Add them up and you have 5x^2 + 4x + 4

OpenStudy (anonymous):

ok so the little triangle thingy is divide

OpenStudy (anonymous):

oh no i understand k give me asec

OpenStudy (anonymous):

now once you have 5xsquared+4x+4 what do you do with the 4?? this is where Im getting all kinds of confused

OpenStudy (ranga):

the little triangle thingy denotes exponent. Since superscripts are not easy on a keyboard we use that ^ instead x^2 = \[x ^{2}\]

OpenStudy (anonymous):

Yes gotcha

OpenStudy (ranga):

OK. Now remember the problem. We converted the left hand side to decimal and we converted the right hand side to decimal. Now just set them equal.

OpenStudy (ranga):

\[5x ^{2} + 4x + 4 = 277\]

OpenStudy (anonymous):

yes ok so we put it together now

OpenStudy (ranga):

Subtract 277 from both sides:

OpenStudy (anonymous):

k

OpenStudy (ranga):

\[5x ^{2} + 4x - 273 = 0\]

OpenStudy (anonymous):

YES i got that far thats where I am stuck since you cant combine x's so do I bring the 4x over???

OpenStudy (ranga):

No. This is called a quadratic equation and there is a formula to find the values of x. Do you know the formula?

OpenStudy (anonymous):

No I wish I did!!!

OpenStudy (anonymous):

I should know it but I dont remember it

OpenStudy (ranga):

oh, you have to look up the formula and become familiar with solving quadratic equations. You will come across it a lot in algebra.

OpenStudy (anonymous):

yes see I was having problems in Algebra so I switched over to Liberal Arts Math and yeah this is not any better plus the book is online and does not go into detail on helping our professor encourages us to search the internet for answers and this does not help me at all..someone that struggles as it is you know what I mean...its a online course if you havent guessed

OpenStudy (ranga):

If the quadratic equation is \[ax ^{2} + bx + c = 0\] then \[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] So to solve any quadratic equation simply compare the general formula with the equation you are trying to solve and find what a, b and c are.

OpenStudy (anonymous):

Oh goodness

OpenStudy (anonymous):

no wonder Im lost

OpenStudy (ranga):

The general quadratic equation is: \[ax ^{2} + bx + c = 0\] The equation that we are trying to solve is: \[5x ^{2} + 4x - 273 = 0\] Compare them term by term and you will recognize a = 5 b = 4 c = -273 Follow?

OpenStudy (anonymous):

is it 6? yes I filled them in sorry I should tell you what Im doing :) Yes I filled them in...

OpenStudy (ranga):

I am getting slightly different answer. First calculate b^2 - 4ac that is inside the square root. a = 5, b = 4, c = -273 4^2 - (4)(5)(-273) 16 + (4)(5)(273) 16 + 5460 5476 Follow

OpenStudy (anonymous):

oH im getting so frustrated worked it like 5 times... ok let me try again

OpenStudy (anonymous):

ok yes then you divide 5476 divided by 10 right?? 2(a) is 10???

OpenStudy (anonymous):

I was actually doing it right!!!!! what you did up top is what I have on my paper

OpenStudy (anonymous):

no its not divided by 10 sorry

OpenStudy (ranga):

First you have to take the square root of 5476

OpenStudy (anonymous):

2,378??

OpenStudy (ranga):

No, square root of 5476 = 74 (you have to use a calculator or you can use Google calculator)

OpenStudy (anonymous):

NO 74

OpenStudy (anonymous):

Oh yeah thats what I put LOl

OpenStudy (anonymous):

then I divide by the 2(a) right

OpenStudy (anonymous):

or is it 10 or 12?? divided into 74 trying to figure

OpenStudy (ranga):

Not yet. A quadratic equation has two solutions. The plus/minus sign you see in the formula means first you have to take it as a plus sign and find x and then you have to take it as a minus sign and find x two get two values for x. \[x = \frac{ -4 \pm 73 }{ (2)(5) } = \frac{ -4 + 73 }{ 10 } or \frac{ -4 - 73}{ 10 }\]

OpenStudy (anonymous):

OH THATS what I did wrong! so which do you use?

OpenStudy (ranga):

OOps. It should be 74 not 73.

OpenStudy (anonymous):

no I got it

OpenStudy (ranga):

\[x = \frac{ -4 + 74 }{ 10} or \frac{ -4 - 74 }{ 10 }\]

OpenStudy (anonymous):

right ...so now I ask which one do I do and when I did both I dont get a even number?? and I have to have a even number for the answer

OpenStudy (ranga):

In this problem you can ignore the second solution because it gives a negative value and for problems involving bases a negative number is not used. So we will use the first solution. -4 + 74 = 70 and when it is divided by 10 you get 7. So x = 7

OpenStudy (anonymous):

OH ok so you want the positive solution then... OH I UNDERSTAND!!!!!! wow that is something that is way over my head!!!!!!!!!!!!!! I have to get some tutoring for all this!!!!!

OpenStudy (ranga):

Quadratic equations and their solutions keep cropping up in Algebra and so it is better to get some practice solving quadratic equations. Also you have to remember the formula for x I gave earlier.

OpenStudy (anonymous):

Right..... yes I agree I need more of these to practice and yes they do pop up I remember some way back in High School geez but have totally slipped my mind now.... I need some extra help.... I sure so appreciate yours!!!!

OpenStudy (anonymous):

Your awesome with math wish I was gifted that way! :)

OpenStudy (ranga):

People have different gifts. There are probably certain things you excel in thatothers dont. So good luck.

OpenStudy (anonymous):

Very true!!! I appreciate it very much and hope you have a wonderful night!!! :)

OpenStudy (ranga):

You too. Bye!

OpenStudy (anonymous):

Bye :)